# Dynamic link switching induces stable synchronized states in sparse networks

### v1

@@ IV Conclusions
In summary, in a double-layer network of FitzHugh-Nagumo (FHN) oscillators, where each layer comprises two species of FHN oscillators oscillating chaotically, we have demonstrated that the switching of sparse inter-layer links can induce asymptotically stable inter-layer synchronized states. More specifically, in a system with $200$ FHN oscillators and a fixed inter-layer coupling intensity, we observe that when the interval between link switches is large, maintaining connections between mirror oscillators in each layer for extended durations is not effective in achieving complete inter-layer synchronization. Conversely, shorter intervals between link switches, where each pair of mirrored oscillators remains connected for briefer periods, promote inter-layer synchronization more effectively. Subsequently, the effectivity of a shorter switching time is demonstrated by estimating the time to complete inter-layer synchronization for different layer sizes.
To estimate the transverse stability of the states of complete inter-layer synchronization, the switching procedure presents challenges due to the discontinuities it may introduce in the system dynamics. To address this issue, we introduced smooth square-wave functions to emulate the link switching. This approach enabled the application of the master stability function formalism, allowing for a stability analysis consistent with the switching procedure in a reduced double-layer network consisting of only two FHN oscillators per layer. Our findings demonstrate that periodically switching the existing links, randomly assigned among the oscillators in a sparse network, can establish stable synchronized states. As open questions for future exploration, we highlight the potential of applying switching connections to control the onset of stable collective states beyond synchronization, such as coherent motion in active systems [39, 40, 41, 42].

### v2

@@ IV Conclusions
In summary, in a double-layer network of FitzHugh–Nagumo (FHN) oscillators, where each layer comprises two species of FHN oscillators exhibiting chaotic dynamics, we have numerically demonstrated that the switching of sparse inter-layer links can induce inter-layer synchronized states. More specifically, for a fixed inter-layer coupling intensity, we observe that when the interval between link switches is large, maintaining connections between mirror oscillators in each layer for extended durations is not effective in achieving complete inter-layer synchronization. Conversely, shorter intervals between link switches, in which each pair of mirrored oscillators remains connected for briefer periods, promote inter-layer synchronization more effectively. Subsequently, the effectiveness of shorter switching times is demonstrated by estimating the time required to achieve complete inter-layer synchronization for different layer sizes.To gain insight into the transverse stability of the observed states of complete inter-layer synchronization induced by intermittently switching links in our double-layers networks, we apply the master stability function (MSF) formalism to a minimal double-layer system containing two oscillators per layer. We first note that the MSF analysis for networks with link switching presents inherent challenges due to the discontinuities that such switching may introduce into the system dynamics. To address this issue, we introduce smooth square-wave functions that emulate the link switching, thereby enabling a well-defined stability analysis. With this approach, we demonstrate that intermittently switching the links can induce stable synchronized states in the minimal system. This strategy captures the fundamental transverse stabilization mechanism induced by link switching.We point out that, in our large double-layer system with sparse inter-layer topologies, synchronization is governed by local transverse modes associated with individual inter-layer connections, rather than by global collective coupling. Fast switching effectively generates a time-averaged stabilizing interaction, and each active inter-layer link locally enforces transverse contraction in the same manner as in the minimal system. Consequently, the stabilization mechanism identified for the minimal system constitutes a fundamental building block of the synchronization process in large sparse double-layer networks, thereby providing a sound justification for extending the minimal-system results on transversal stability to the large-scale regime considered here.We emphasize that spontaneous intra-layer synchronization would reduce the effective layer dimensionality, thereby facilitating the onset of inter-layer synchronization through the switching strategy. Accordingly, the use of two oscillator populations within each layer represents a first step toward introducing heterogeneities that suppress spontaneous intra-layer synchronization. We further note that stronger dynamical heterogeneities at the oscillator level, or the introduction of topological heterogeneity such as random intra-layer connections, would pose additional difficulties for the establishment of inter-layer synchronization via link switching. Therefore, the investigation of different types of intra-layer heterogeneities is left for future work. As additional open questions for future exploration, we highlight the potential of applying switching connections to control the onset of stable collective states beyond synchronization, such as coherent motion in active systems [11, 24, 39, 20], polarized opinion patterns in social groups [1, 27], and spreading patterns of epidemics [26, 10].